# Zero-TIG: Temporal Consistency-Aware Zero-Shot Illumination-Guided Low-light Video Enhancement

### v1

@@ Abstract
Low-light and underwater videos suffer from poor visibility, low contrast, and high noise, necessitating enhancements in visual quality. However, existing approaches typically rely on paired ground truth, which limits their practicality and often fails to maintain temporal consistency. To overcome these obstacles, this paper introduces a novel zero-shot learning approach named Zero-TIG, leveraging the Retinex theory and optical flow techniques. The proposed network consists of an enhancement module and a temporal feedback module. The enhancement module comprises three subnetworks: low-light image denoising, illumination estimation, and reflection denoising. The temporal enhancement module ensures temporal consistency by incorporating histogram equalization, optical flow computation, and image warping to align the enhanced previous frame with the current frame, thereby maintaining continuity. Additionally, we address color distortion in underwater data by adaptively balancing RGB channels. The experimental results demonstrate that our method achieves low-light video enhancement without the need for paired training data, making it a promising and applicable method for real-world scenario enhancement.
@@ IV-B Performance for low-light video enhancement
Table I presents the results, evaluated using well-known metrics: PSNR, SSIM, and LPIPS, as the BVI-RLV dataset provides normal light references. Also, since brightness can be subjective, we included results after applying histogram matching (HM) to the references to remove brightness differences, allowing us to evaluate the denoising performance of each method more accurately. These results demonstrate that our Zero-TIG outperforms most methods and is relatively comparable to EnlightenGAN. Although EnlightenGAN achieves the highest score on the LPIPS metric, it is prone to generating hallucinated color artifacts in images, as subjectively evaluated and illustrated in Fig. 3. Compared to Zero-IG, our method demonstrates superior denoising performance. Although supervised methods achieve higher quantitative metrics, our visual results are comparable to, and in some cases even surpass, those of supervised approaches, outperforming BVI-CDM in terms of visual quality.
@@ IV-C Performance of low-light underwater video enhancement
Our results demonstrate superior contrast compared to Zero-IG, with clearer edges of coral and sands in the background, as illustrated in Fig. 5. Quantitatively, our method achieves higher scores across all non-reference metrics. These improvements confirm that our approach effectively addresses distortions in underwater videos and achieves adaptive white balance, enhancing both visual quality and metric performance.

### v2

@@ Abstract
Low-light and underwater videos suffer from poor visibility, low contrast, and high noise, necessitating enhancements in visual quality. However, existing approaches typically rely on paired ground truth, which limits their practicality and often fails to maintain temporal consistency. To overcome these obstacles, this paper introduces a novel zero-shot learning approach named Zero-TIG, leveraging the Retinex theory and optical flow techniques. The proposed network consists of an enhancement module and a temporal feedback module. The enhancement module comprises three subnetworks: low-light image denoising, illumination estimation, and reflection denoising. The temporal enhancement module ensures temporal consistency by incorporating histogram equalization, optical flow computation, and image warping to align the enhanced previous frame with the current frame, thereby maintaining continuity. Additionally, we address color distortion in underwater data by adaptively balancing RGB channels. The experimental results demonstrate that our method achieves low-light video enhancement without the need for paired training data, making it a promising and applicable method for real-world scenario enhancement. Code is available at https://github.com/liyinibristol/Zero-TIG.
@@ IV-B Performance for low-light video enhancement
Table I presents the results, evaluated using well-known metrics: PSNR, SSIM, and LPIPS, as the BVI-RLV dataset provides normal light references. Also, since brightness can be subjective, we included results after applying histogram matching (HM) to the references to remove brightness differences, allowing us to evaluate the denoising performance of each method more accurately. These results demonstrate that our Zero-TIG outperforms most methods and is relatively comparable to Zero-IG. The EnlightenGAN is prone to generating hallucinated color artifacts in images, as subjectively evaluated and illustrated in Fig. 3. Compared to Zero-IG, our method demonstrates superior denoising performance. Although supervised methods achieve higher quantitative metrics, our visual results are comparable to, and in some cases even surpass, those of supervised approaches, outperforming BVI-CDM in terms of visual quality.
@@ IV-C Performance of low-light underwater video enhancement
Our results demonstrate superior contrast compared to Zero-IG, with clearer edges of coral and sand in the background, as illustrated in Fig. 5. Quantitatively, our method achieves higher scores across all non-reference metrics. These improvements confirm that our approach effectively addresses distortions in underwater videos and achieves adaptive white balance, enhancing both visual quality and metric performance.